OpenStudy (kydawg):

Verify the identity. cot (x-π/2) = -tan x

satellite73 (satellite73):

depends on what you get to use the cofunction identity tells you \[\cot(\frac{\pi}{2}-x)=\tan(x)\] and cotangent is an odd function so \[\cos(x-\frac{\pi}{2})=-\tan(x)\]

OpenStudy (kydawg):

I don't get it.

sam (.sam.):

satellite73 explained it already, cotangent is an odd function so, \[\cot(-y)=-\tan(y)\]

jhonyy9 (jhonyy9):

@satellite73 sorry here you missed the cot and wrote cos << and cotangent is an odd function so cos(x−π2)=−tan(x) >> - this above wann being cot(x-pi/2)=-tanx

OpenStudy (kydawg):

so cot(π/2-x) = tanx?

satellite73 (satellite73):

yes

satellite73 (satellite73):

that is a "cofunction" identity like \[\sin(\frac{\pi}{2}-x)=\cos(x)\] and \[\csc(\frac{\pi}{2}-x)=\sec(x)\] if you draw a right triangle you will see why